OpenStudy (anonymous):

(7,14) and (8,16) do n e body know how to to get the slope intercept and point slope form of this

OpenStudy (anonymous):

point slope formula is y-y=m(x-x) slope intercept form i belive is y=mx+b

OpenStudy (anonymous):

to find slope which is m rise / run y-y/m-m

OpenStudy (anonymous):

can u help me answer it

OpenStudy (anonymous):

just plug in the numbers try it out and ill tell you if your doing it right

OpenStudy (anonymous):

you know which numbers are x and which are y right?

OpenStudy (anonymous):

i ment y-y/x-x btw not m-m sorry

OpenStudy (anonymous):

I DONT THINK I AM

OpenStudy (anonymous):